OpenStudy (anonymous):

p(x) = 2x^3-x^2-3x+1 (3) (-2)

OpenStudy (gw2011):

If you are trying to find the value of the equation for 3 and for -2, then you would solve it the following way: p(3)=2(3)^3-(3)^2-3(3)+1 p(3)=54-9-9+1 p(3)=37 p(-2)=2(-2)^3-(-2)^2-3(-2)+1 p(-2)=-16-4+6+1 p(-2)=-13